What is the relation between productivity and cost? Describe the association using equation for MC & MP and AP & AC


1
Expert's answer
2021-06-15T11:16:43-0400

In the United States, productivity is a metric that quantifies labor efficiency in the production of products and services. The cost indicator in the US economy measures the unit labor expenses of producing each unit of output.

"AP = \frac{TP}{Q}"

"MP = TPn - TP_{n-1}"

"TC = TFC + TVC"

"AFC = \frac{TFC}{Q}"

"AC = \frac{TC}{Q}"

"MC = TVC_n - TVC_{n-1}"

a) Relationship between AVC and MC

- AVC falls till when "AVC > MC"

- AVC becomes minimum when "AVC = MC"

- AVC starts rising when AVC < MC

- MInimum point of AVC is to the right of the minimum of MC

b)  Relationship between AVC and MC

- ATC falls till when ATC > MC

- ATC becomes minimum when ATC = MC

- ATC starts rising when "ATC < MC"

- MInimum point of ATC is to the right of a minimum of MC

c) Relationship between AP and AVC

AVC is the reciprocal of AP

- When AP increases in the short run AVC falls

- When AP falls in the short run AVC rises

Relationship between MP and MC

MC is the reciprocal of MP

- When MP increases in the short run MC falls

- When MP falls in the short run MC rises